# On the diameter of intersection graphs of finite groups

### 07 Mar 2024 v1

@@ Abstract.
Let $G$ be a finite group. The intersection graph $\Delta_{G}$ of $G$ is a simple graph with vertices the non-trivial proper subgroups of $G$, and an edge between two vertices if the corresponding subgroups intersect non-trivially. These graphs were introduced by Csákány and Pollák [6] in 1969, influenced by earlier work by Bosák [1] on analogous graphs for semigroups. In their paper, Csákány and Pollák prove that a finite non-simple group with disconnected intersection graph either has trivial centre and every proper subgroup is abelian, or it is a direct product of two cyclic groups of prime order. They further show that the intersection graph of every other non-simple finite group has $\mathrm{diam}(\Delta_{G})\leq 4$ with equality only if $G$ is an almost simple group of the form $G=S\rtimes C_{p}$ for some non-abelian finite simple group and prime $p$. Freedman further showed that in fact, $p$ must be an odd prime [8]. Despite this reduction, no example of a non-simple group with intersection graph of diameter 4 has been found to date. We change this, by presenting a group whose intersection graph has diameter 4, as well as providing some necessary conditions for any further examples.
@@
    $G_{0}=\mathrm{PSL}_{n}(q)$ with $n=p$ prime, and $n\mid q-1$,
@@
    $G_{0}=\mathrm{PSU}_{n}(q)$ with $p\mid q+1$, or
@@ 1. Non-simple groups with connected intersection graphs
We first begin with some preliminary results.
@@
We start by describing the groups $G$ such that $\Delta_{G}$ has diameter $2$
@@
Let $S$ be a finite non-abelian simple group of Lie type. Then any non-diagonal outer automorphism of odd order of $S$ normalizes a parabolic maximal subgroup of $S$. Moreover, any two of these normalise parabolic maximal subgroups that are conjugate.
@@
First note that $\mathrm{Aut}(S)$ is a split extension of $\mathrm{InnDiag}(S)$ by a group $\Phi_{S}\Gamma_{S},$ where $\Phi_{S}$ and $\Gamma_{S}$ contain field and graph automorphisms, respectively. By [11, Thm. 2.5.12, Defn. 2.5.13], any non-diagonal outer automorphism $g$ of $S$ of odd prime order is a field automorphism, or $S\cong\mathrm{P}\Omega^{+}_{8}(q)$ and $g$ is a graph or graph-field automorphism. In all cases, $g$ is conjugate to an element of $\Phi_{S}\Gamma_{S}$.
@@
If $g_{1}$ and $g_{2}$ are field automorphisms, then, in particular, they are each conjugate to an element of $\Phi_{S}$ by [11, Defn. 2.5.13]. Now $\Phi_{S}$ normalises all parabolic subgroups of $S$ by [11, §2.6], so the claim follows. If $g_{1}$ and $g_{2}$ are graph or graph-field automorphisms, then $S\cong\mathrm{P}\Omega^{+}_{8}(q),$ and $S$ contains a parabolic maximal subgroup $P$ normalized by $\Gamma_{S}$ by [3, Table 8.50]. Hence $P$ is normalized by $\Phi_{S}\Gamma_{S},$ and as all non-diagonal automorphisms are conjugate to an element of $\Phi_{S}\Gamma_{S},$ $g_{1}$ and $g_{2}$ both normalize a conjugate of $P.$
@@
By [11, Prop. 2.6.9] (see also [3, Chap 8.] and [13, Tables 3.5A–G]), an analogue of Lemma 5 holds for diagonal automorphisms that are conjugate to the so-called “standard” diagonal automorphisms (cf. [3, §1.7.1]). However, not every diagonal automorphism is conjugate to a standard one, and therefore not always contained in a parabolic subgroup, as we shall see in the proof of Theorem 1.
@@ 1.1. Proof of Theorem 1
Our second case is when $S=PSL_{n}(q)$ such that $n$ is composite. Then by [4, Prop. 3.2.2] $A$ and $B$ both fix a $p$-dimensional subspace of the natural module of $PSL_{n}(q).$ In this case are exactly the stabilizers of $p$-dimensional subspaces by [17, Section 3.3.3], so they are contained in conjugate parabolic subgroups.
By [15, Lemma 4.1.5], [10, Corollary 1.3] and [5, Thm 2.8.7], the intersection of two conjugate parabolics is always non-trivial. Hence in both of the above cases there is a path of length at most $3$ between any two prime order subgroups of $G,$ so $\mathrm{diam}(\Delta_{G})\leq 3.$
@@ 2.1. Proof of Theorem 2
For the remainder of the proof, we therefore suppose either $n=19$, or $n\geq 29$. Let $A=\langle g_{a}\rangle$ and $B=\langle g_{b}\rangle$, where $g_{a}=(1,2,\dots,n-2,n-1,n)$ and $g_{b}=g_{a}^{(n-1,n)}=(1,2,\dots,n-2,n,n-1)$. Every maximal subgroup of $A_{n}$ containing $g_{a}$ or $g_{b}$ must be transitive of prime degree. Thus, by Burnside’s Theorem [7, p.99], the only maximal subgroup of $G$ containing $A$ is $N_{G}(A)\cong n:\frac{n-1}{2}$, and similarly, the only maximal subgroup of $G$ containing $B$ is $N_{G}(B)\cong n:\frac{n-1}{2}$. Suppose $1\neq h\in N_{G}(A)\cap N_{G}(B)$. Note that $h$ has order dividing $\frac{n-1}{2}$, since $A$ and $B$ are distinct. Moreover, $h$ preserves a partition of $\{1,\dots n\}$ into two parts of size $\frac{n-1}{2}$, and one part of size 1. In particular, $h$ fixes exactly one point. Relabelling via a conjugate of $g_{a}$ if necessary, we may suppose that $h$ fixes 1, so now $g_{a}=(1,2,\dots,n-2,n-1,n)$ and $g_{b}=g_{a}^{(k,k+1)}$ with $11\leq k\leq n-$ 1. Hence $g_{a}^{h}=(1,h(2),h(3),\dots,h(n))$ and the $i^{th}$ position of $g_{b}^{h}$ is similarly $h(i)$ unless $i=k$ or $k+1,$ in which case it is $h(k+1)$ and $h(k),$ respectively.
@@
Note that $g_{a}$ and $g_{b}$ agree on every element except for those with entries $k$, $k+1$. Similarly, the powers $g_{a}^{i}$ and $g_{b}^{i}$ also agree on every element except those with entries $k$ or $k+1$. In particular, if $k$ is in position $l^{\prime}$ in $g_{a}^{i},$ then $k+1$ is in position $l^{\prime}$ for $g_{b}^{i},$ so there exists $\ell^{\prime}\in\{1,\dots,n\}$ such that $g_{a}^{i}(\ell^{\prime})=k$ and $g_{b}^{i}(\ell^{\prime})=k+1$. We now set out to show that $\ell^{\prime}\not\in\{1,k,k+1\}$. First note that $\ell^{\prime}\neq 1$ by assumption. If $\ell^{\prime}=k$, then since $n$ is prime, $A$ acts sharply transitively on $\{1,\dots,n\}$, $g_{a}^{i}(k)=k$ implies $i=1$. This is a contradiction since $h$ does not centralise $g_{a}$. If instead $\ell^{\prime}=k+1$, then a similar argument for $B$ gives another contradiction. Therefore, $\ell^{\prime}\not\in\{1,k,k+1\}$. Hence,
$k=(\ell^{\prime})g_{a}^{i}=(\ell^{\prime})g_{a}^{h}=(\ell^{\prime})h$
and also
$k+1=(\ell^{\prime})g_{b}^{i}=(\ell^{\prime})g_{b}^{h}=(\ell^{\prime})h,$
a contradiction. It follows $h$ does not exist, $N_{G}(A)\cap N_{G}(B)=\{1\}$ and therefore $P$ and $Q$ are at distance at least 4 in $\Delta_{G}$ and so $\mathrm{diam}(\Delta_{G})=4$.

### 20 Mar 2024 v2

@@ Abstract.
Let $G$ be a finite group. The intersection graph $\Delta_{G}$ of $G$ is a simple graph with vertices the non-trivial proper subgroups of $G$, and an edge between two vertices if the corresponding subgroups intersect non-trivially. These graphs were introduced by Csákány and Pollák [6] in 1969, influenced by earlier work by Bosák [1] on analogous graphs for semigroups. In their paper, Csákány and Pollák prove that a finite non-simple group with disconnected intersection graph either has trivial centre and every proper subgroup is abelian, or it is a direct product of two cyclic groups of prime order. They further show that the intersection graph of every other non-simple finite group has $\mathrm{diam}(\Delta_{G})\leq 4$ with equality only if $G$ is an almost simple group of the form $G=S\rtimes C_{p}$ for some non-abelian finite simple group and prime $p$. Freedman further showed that in fact, $p$ must be an odd prime [9, p.83]. Despite this reduction, no example of a non-simple group with intersection graph of diameter 4 has been found to date. We change this, by presenting a group whose intersection graph has diameter 4, as well as providing some necessary conditions for any further examples.
@@
    $G_{0}=\mathrm{PSL}_{n}(q)$ with $n=p$ prime, and $n\mid q-1$;
@@
    $G_{0}=\mathrm{PSU}_{n}(q)$ with $p\mid(n,q+1)$, and either $n=p$ prime, or $n=2p$;
@@ 1. Non-simple groups with connected intersection graphs
We first begin with some preliminary results. Note that this result is also in [6].
@@
We start by describing the groups $G$ such that $\Delta_{G}$ has diameter $2$ Note that one direction of this result is also in [17].
@@
Let $S$ be a finite non-abelian simple group of Lie type. Then any non-diagonal outer automorphism of odd prime order of $S$ normalizes a parabolic maximal subgroup of $S$. Moreover, any two of these automorphisms normalise parabolic maximal subgroups that are conjugate.
@@
First note that $\mathrm{Aut}(S)$ is a split extension of $\mathrm{InnDiag}(S)$ by a group $\Phi_{S}\Gamma_{S},$ where $\Phi_{S}$ and $\Gamma_{S}$ contain field and graph automorphisms, respectively. By [12, Thm. 2.5.12, Defn. 2.5.13], any non-diagonal outer automorphism $g$ of $S$ of odd prime order is a field automorphism, or $S\cong\mathrm{P}\Omega^{+}_{8}(q)$ or ${}^{3}D_{4}(q)$ and $g$ is a graph or graph-field automorphism. In all cases, $g$ is conjugate to an element of $\Phi_{S}\Gamma_{S}$.
@@
If $g_{1}$ and $g_{2}$ are field automorphisms, then, in particular, they are each conjugate to an element of $\Phi_{S}$ by [12, Defn. 2.5.13]. Now $\Phi_{S}$ normalises all parabolic subgroups of $S$ by [12, §2.6], so the claim follows. If $g_{1}$ and $g_{2}$ are graph or graph-field automorphisms, then $S\cong\mathrm{P}\Omega^{+}_{8}(q)$ or ${}^{3}D_{4}(q)$ and the parabolic maximal subgroup $P_{2}$ of $S$ is normalized by $\Gamma_{S}.$ Hence $P_{2}$ is normalized by $\Phi_{S}\Gamma_{S},$ and as all non-diagonal automorphisms are conjugate to an element of $\Phi_{S}\Gamma_{S},$ $g_{1}$ and $g_{2}$ both normalize a conjugate of $P_{2},$ as required. ∎
@@
By [12, Prop. 2.6.9] (see also [3, Chap 8.] and [14, Tables 3.5A–G]), an analogue of Lemma 5 holds for diagonal automorphisms that are conjugate to the so-called “standard” diagonal automorphisms (cf. [3, §1.7.1]). However, not every diagonal automorphism is conjugate to a standard one, and therefore not always contained in a parabolic subgroup, as we shall see in the proof of Theorem 1.
@@ 1.1. Proof of Theorem 1
Our second case is when $S=PSL_{n}(q)$ such that $n$ is composite and $A$ and $B$ are generated by diagonal automorphisms. Then by [4, Prop. 3.2.2] $A$ and $B$ both fix a $p$-dimensional subspace of the natural module of $PSL_{n}(q).$ In this case are exactly the stabilizers of $p$-dimensional subspaces by [19, Section 3.3.3], so they are contained in conjugate parabolic subgroups.
By [16, Lemma 4.1.5], [11, Corollary 1.3] and [5, Thm 2.8.7], the intersection of two conjugate parabolics is always non-trivial. Hence in both of the above cases there is a path of length at most $3$ between any two prime order subgroups of $G,$ so $\mathrm{diam}(\Delta_{G})\leq 3.$
Our third case is $S=\mathrm{PSU}_{n}(q)$ with $n$ composite and $A$, $B$ generated by diagonal automorphisms of order $p$ dividing $(n,q+1)$. By [4, Prop. 3.3.3], the preimage of each of these elements in $\mathrm{GU}_{n}(q)$ each fixes an orthogonal decomposition of the natural module into non-degenerate subspaces. Hence, each is contained in a maximal subgroup $M$ of $G$ whose preimage in $\mathrm{GU}_{n}(q)$ preserves an orthogonal decomposition of the natural module into two complementary non-degenerate subspaces of dimension $m$ and $n-m$ for some $21\leq m\leq n/$ 2 respectively. By [14, Tables 2.1.C, 2.1.D, Lemma 4.1.1], we deduce that if $m<n/2$, then $|M|>|G|^{1/2}$. Therefore, if $A$ and $B$ are generated by elements that fix a decomposition of the natural module into a pair of non-degenerate subspaces of unequal dimension, then the maximal subgroups $M_{1}\geq A$, $M_{2}\geq B$ of $G$ fixing these decompositions must have $|G|<|M_{1}||M_{2}|=|M_{1}M_{2}||M_{1}\cap M_{2}|=|G||M_{1}\cap M_{2}|$. This implies $|M_{1}\cap M_{2}|>1$, so $A$, $B$ are at distance 3 in $\Delta_{G}$.
@@
Remark 1.2.
Note that the only cases remaining when $S=\mathrm{PSU}_{n}(q)$ in the proof of Theorem 1 are those where $g$ is a diagonal automorphism of prime order $p$ whose preimage in $\mathrm{GU}_{n}(q)$ does not preserve a decomposition of the natural module into two non-degenerate subspaces of unequal dimension. By [4, Prop. 3.3.3], this implies that either $n$ is prime, or $n=2p$.
@@ 2.1. Proof of Theorem 2
For the remainder of the proof, we therefore suppose either $n=19$, or $n\geq 29$. Let $A=\langle g_{a}\rangle$ and $B=\langle g_{b}\rangle$, where $g_{a}=(1,2,\dots,n-2,n-1,n)$ and $g_{b}=g_{a}^{(n-1,n)}=(1,2,\dots,n-2,n,n-1)$. Every maximal subgroup of $A_{n}$ containing $g_{a}$ or $g_{b}$ must be transitive of prime degree. Thus, by the classification of transitive groups of prime degree [7, p.99 Claims (i)-(v)], the only maximal subgroup of $G$ containing $A$ is $N_{G}(A)\cong n:\frac{n-1}{2}$, and similarly, the only maximal subgroup of $G$ containing $B$ is $N_{G}(B)\cong n:\frac{n-1}{2}$. Suppose $1\neq h\in N_{G}(A)\cap N_{G}(B)$. Note that $h$ has order dividing $\frac{n-1}{2}$, since $A$ and $B$ are distinct. Moreover, $h$ preserves a partition of $\{1,\dots n\}$ into two parts of size $\frac{n-1}{2}$, and one part of size 1. In particular, $h$ fixes exactly one point. Relabelling via a conjugate of $g_{a}$ if necessary, we may suppose that $h$ fixes 1, so now $g_{a}=(1,2,\dots,n-2,n-1,n)$ and $g_{b}=g_{a}^{(k,k+1)}$ with $11\leq k\leq n-$ 1. Hence $g_{a}^{h}=(1,(2)h,(3)h,\dots,(n)h$ and the $i^{th}$ position of $g_{b}^{h}$ is similarly $(i)h$ unless $i=k$ or $k+1,$ in which case it is $(k+1)h$ and $(k)h,$ respectively.
@@
Since $g_{b}=g_{a}^{(k,k+1)},$ also $g_{b}^{i}=(g_{a}^{i})^{(k,k+1)}=(g_{a}^{h})^{(k,k+1)}.$ Without loss of generality we can assume that there is >′′$\ell^{\prime}>\ell^{\prime\prime}>1$ such that $(\ell^{\prime})h=k$ and (′′$(\ell^{\prime\prime})h=k+1.$ Note that unless $i=1,$′$\ell^{\prime}-\ell^{\prime\prime}\geq 2,$ so at least one of them is not $k$ or $k+1.$ Now $g_{b}^{h}=(1,\dots,(\ell^{\prime}-1)h,(\ell^{\prime})h,(\ell^{\prime}+1)h,\dots,(\ell^{\prime\prime}-1)h,(\ell^{\prime\prime})h,(\ell^{\prime\prime}+1)h,\dots)$ and also, $g_{b}^{h}=(g_{a}^{h})^{(k,k+1)}=1,\dots,(\ell^{\prime}-1)h,(\ell^{\prime\prime})h,(\ell^{\prime}+1)h,\dots,(\ell^{\prime\prime}-1)h,(\ell^{\prime})h,(\ell^{\prime\prime}+1)h,\dots).$ Hence $(\ell^{\prime})h=(\ell^{\prime\prime})h,$ which is a contradiction.
It follows $h$ does not exist, $N_{G}(A)\cap N_{G}(B)=\{1\}$ and therefore $P$ and $Q$ are at distance at least 4 in $\Delta_{G}$ and so $\mathrm{diam}(\Delta_{G})=4$.